when solving the system of equations, 3x+y=14
which expression could be substituted x+4y=3
for y in the second equation?
a.3-4y
b.3-x/4
c.14-y/3
d.14-3x

Answer by 303795(562)

3x + y = 14
Rearrange the equation to get y on its own.
Subtract 3x from each side
3x + y - 3x = 14 - 3x
y = 14 - 3x
This can now be substituted in the second equation.